OpenStudy (anonymous):

find volume of solid that results when th region enclosed by hte curves is revolved about the y-axis ... x=2, x=1/y, y=3

OpenStudy (anonymous):

Evaluate the integral\[\int\limits_{0.5}^{3} \pi \frac{ 1 }{ y^2 } dy + \int\limits_{0}^{0.5} \pi/4 dy\] There you have your answer. Why the integral is split? Graph the function and you'll get it

OpenStudy (anonymous):

thanks